OpenStudy (anonymous):

If the midpoint is the point (2,4) and one endpoint is (6,12) find the other endpoint

OpenStudy (anonymous):

Let the other endpoint be (h,k) Therefore, (6 + h)/2 = 2 6 + h = 4 h = -2 (12 + k) / 2 = 4 12 + k = 8 k = -4 The other endpoint is (-2, -4)